OpenStudy (anonymous):

Find the domain of y = 2x + 1 if the range is {9, 13, 19, 23} Select one: a. {4, 5, 6, 7} b. {4, 6, 9, 11} c. {19, 27, 39, 47} d. {5, 7, 10, 12}

OpenStudy (amistre64):

solve for x = elements of the domain

OpenStudy (amistre64):

hmm, i see its earlier and i read it wrong

OpenStudy (amistre64):

you can rearrange it as x=my+d if you want and run it

OpenStudy (amistre64):

or, use the options as your x values to see what produces the given range

OpenStudy (anonymous):

y = 2 * 9 + 1 = 10 y = 2 * 13 + 1 = 27 y = 2 * 19 + 1 = 39

OpenStudy (anonymous):

SO C

OpenStudy (amistre64):

no c, that was a bait .... i feel for it too

OpenStudy (amistre64):

which option produces the given range

OpenStudy (anonymous):

huh ok..... I don't get it

OpenStudy (amistre64):

y = 2x + 1, they give you the y values {9, 13, 19, 23} 9 = 2x + 1 13 = 2x + 1 19 = 2x + 1 23 = 2x + 1 and ask you for the set of x values that work

OpenStudy (anonymous):

{4, 5, 6, 7} b. {4, 6, 9, 11} c. {19, 27, 39, 47} d. {5, 7, 10, 12}

OpenStudy (anonymous):

fill in x?

OpenStudy (amistre64):

correct, fill in x

OpenStudy (anonymous):

2 * 4 + 1 = 9 2 * 6 + 1 =13 2 * 9 + 1 = 19 2 * 11 + 1 = 23 B?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

Thanks :D

OpenStudy (amistre64):

youre welcome